Form a polynomial f(x) with real coefficients having the given degree and zeros.
Degree: 5; zeros: 2, -3i, and 4 - i

Answer by ewatrrr(24785)

(x-2)(x^2 + 9)[x-(4-i)][x-(4+i)] = (x-2)(x^2+9)[(x-4)^2 - i^2]
(x-2)(x^2+9)[(x-4)^2 - i^2] = (x-2)(x^2+9)(x^2 - 8x + 16 + 1}
(x-2)(x^2+9)(x^2 - 8x + 17}
Check Arithmetic
Multiply to finish